Power in the rate at which work is done or energy is transformed. That is,

Power= Energy/Time

Use the energy found above and the time measured during the demonstration to calculate the power developed in the eruption.

Starting Mass 4.6 lbs

Max height 41 inches

Time of Eruption 3.35 sec

Final Mass after Eruption: 1.7 lbs

Amount of Mass that left the bottle during Eruption 2.9lbs

Please help me answer the above question with the results from the experiment.
Answer #1

Energy converted into potential energy = m×g×h = 1.315 × 9.8 × 41× 0.0254 J = 13.42 J

Power developed = Energy /Time = 13.42/3.35 = 4 W